# Meta-Statistical Learning: Supervised Learning of Statistical Inference

### v1

@@ 1 Introduction
Contributions. In this work: (i) We introduce the meta-statistical framework and establish its connections to related paradigms such as meta-learning and amortized inference (Section 2). (ii) We evaluate various meta-statistical architectures, demonstrating that dataset encoders based on Set Transformer (Lee et al., 2019) variants achieve strong performance and generalization while avoiding the quadratic computational cost of standard attention mechanisms (Section 4). (iii) We showcase the versatility of the framework through three statistical inference problems: estimating standard deviation, conducting normality tests, and estimating mutual information. Our results highlight the robustness and generalization capabilities of meta-statistical models across diverse data distributions and, especially in low-sample-size settings (Section 5). (iv) We release for practitioners meta-statistical normality test and MI estimator, strong in low-sample regimes.
@@ 2.4 Related Work
Relationship to Meta-Learning. Meta-learning, or learning to learn, is a paradigm focused on generalizing across tasks drawn from different distributions (Schmidhuber et al., 1996; Hospedales et al., 2021; Huisman et al., 2021). Meta-learning seeks to acquire transferable meta-knowledge, enabling rapid adaptation to new tasks (Schmidhuber, 1987; Thrun, 1998; Schmidhuber, 1993; Vanschoren, 2019). A broad range of approaches exist (Vinyals et al., 2016; Santoro et al., 2016; Finn et al., 2017; Snell et al., 2017), some emphasizing dataset-level processing to extract useful representations (Mishra et al., 2017; Ravi and Larochelle, 2017; Munkhdalai and Yu, 2017; Shyam et al., 2017). This is particularly relevant in few-shot learning (Finn et al., 2017; Snell et al., 2017; Wang et al., 2023; Wu et al., 2020; Rivolli et al., 2022). Notably, neural processes represent a class of meta-learners that use a meta-distribution over functions, adapting their prior to new datasets using observed input-output pairs (Garnelo et al., 2018b, a; Kim et al., 2019). Meta-statistical learning shares conceptual similarities with meta-learning, as both focus on generalization across distributions. However, while the target of meta-learning remains instance-level predictions, meta-statistical learning emphasizes distributional properties. These paradigms are complementary: insights from dataset-level analysis can directly improve generalization in meta-learning (Jomaa et al., 2021; Kotlar et al., 2021).
@@ 6 Discussion
Interpretability is another challenge. The precise algorithm computed at inference to perform the statistical inference becomes unknown and difficult to interpret (Teney et al., 2022). Also, like LLMs, meta-statistical models could exhibit unexpected failure cases and lack strict guarantees of validity. However, they also offer a promising testbed for mechanistic interpretability research: they process structured numerical inputs without tokenization, operate in a single forward pass, and construct mathematical representations rather than linguistically ambiguous ones.

### v2

@@ 1 Introduction
This perspective builds on a growing body of work that seeks to replace or assist hand-crafted statistical procedures with trained models. Approaches such as neural processes (Garnelo et al., 2018b, a; Gordon et al., 2020; Markou et al., 2022; Huang et al., 2023; Bruinsma et al., 2023; Chang et al., 2025), simulation-based inference (Papamakarios et al., 2019; Cranmer et al., 2020), and amortized Bayesian inference (Chan et al., 2018; Chen et al., 2023, 2020; Elsemüller et al., 2024; Radev et al., 2020; Avecilla et al., 2022; Gloeckler et al., 2024) have demonstrated the feasibility of learning statistical inference procedures directly from data. Meta-statistical learning focuses on the core challenge of learning statistical estimators. This enables the development of models that adapt their estimation strategies to new data distributions in a fully data-driven manner.
@@
Contributions. In this work: (i) We introduce the meta-statistical framework and establish its connections to related paradigms such as meta-learning and amortized inference (Section 2). (ii) We evaluate various meta-statistical architectures, demonstrating that dataset encoders based on Set Transformer (Lee et al., 2019) variants achieve strong performance and generalization while avoiding the quadratic computational cost of standard attention mechanisms (Section 4). (iii) We showcase the versatility of the framework through three statistical inference problems: estimating standard deviation, conducting normality tests, and estimating mutual information. Our results highlight the robustness and generalization capabilities of meta-statistical models across diverse data distributions and, especially in low-sample-size settings (Section 5).
@@ 2.4 Related Work
Machine Learning for Statistical Inference. Our work aligns with the broader research direction on neural processes (Garnelo et al., 2018b, a; Kim et al., 2019; Gordon et al., 2020; Markou et al., 2022; Huang et al., 2023; Bruinsma et al., 2023). Neural processes can predict latent variables of interest from datasets (Chang et al., 2025) by leveraging transformers (Nguyen and Grover, 2022) and Deep Sets (Zaheer et al., 2017) to enforce permutation invariance (Bloem-Reddy and Teh, 2020). A related approach, known as prior-fitted networks, has demonstrated that transformers can be effectively repurposed for Bayesian inference (Müller et al., 2022) and optimization tasks (Müller et al., 2023).
Additionally, there is growing interest in using trained models to assist in statistical inference (Angelopoulos et al., 2023) and optimization (Lueckmann et al., 2017; Liu et al., 2020; Simpson et al., 2021; Amos, 2023). In particular, simulation-based inference benefits from neural simulations (Papamakarios et al., 2019; Cranmer et al., 2020) and amortized Bayesian inference (Gonçalves et al., 2020; Elsemüller et al., 2024; Radev et al., 2020; Avecilla et al., 2022; Gloeckler et al., 2024). Amortized Bayesian inference typically replaces probabilistic inference with a neural network prediction task (Chan et al., 2018; Chen et al., 2023, 2020). These previous work illustrate the feasibility of learning distribution-relevant parameters via maximum likelihood using permutation-invariant dataset representations. In this work, we identify the emerging theme: translate complex statistical inference problems and into the powerful and flexible framework of supervised learning. We then undertake a study of this paradigm from the ground up and investigate parameter efficient dataset encoders like the Set Transformer (Lee et al., 2019; Zhang et al., 2022).
Relationship to Meta-Learning. Meta-learning, or learning to learn, is a paradigm focused on generalizing across tasks drawn from different distributions (Schmidhuber et al., 1996; Hospedales et al., 2021; Huisman et al., 2021). Meta-learning seeks to acquire transferable meta-knowledge, enabling rapid adaptation to new tasks (Schmidhuber, 1987; Thrun, 1998; Schmidhuber, 1993; Vanschoren, 2019). A broad range of approaches exist (Vinyals et al., 2016; Santoro et al., 2016; Finn et al., 2017; Snell et al., 2017), some emphasizing dataset-level processing to extract useful representations (Mishra et al., 2017; Ravi and Larochelle, 2017; Munkhdalai and Yu, 2017; Shyam et al., 2017). This is particularly relevant in few-shot learning (Finn et al., 2017; Snell et al., 2017; Wang et al., 2023; Wu et al., 2020; Rivolli et al., 2022). Notably, neural processes represent a class of meta-learners that use a meta-distribution over functions, adapting their prior to new datasets using observed input-output pairs (Garnelo et al., 2018b, a; Kim et al., 2019). Meta-statistical learning shares conceptual similarities with meta-learning, as both focus on generalization across distributions. However, while the target of meta-learning remains instance-level predictions, meta-statistical learning emphasizes distributional properties. These paradigms are complementary: insights from dataset-level analysis can directly improve generalization in meta-learning (Jomaa et al., 2021; Kotlar et al., 2021; Kobalczyk and van der Schaar, 2025).
@@ 6 Discussion
Statistical inference is fundamentally constrained by irreducible errors arising from finite sample sizes, imposing inherent limits on any estimator’s performance (Casella and Berger, 2024). Consequently, no meta-statistical estimators can surpass these fundamental limits. However, the approach provides a flexible framework for finding estimators with a desired bias-variance trade-off by modifying the loss function. Moreover, the approach allows for the incorporation of complex prior information through the choice of meta-distribution, effectively guiding the estimator’s behavior in a principled manner.
@@
Interpretability is another challenge. The precise algorithm computed at inference to perform the statistical inference becomes unknown and difficult to interpret (Molnar, 2022; Carvalho et al., 2019; Teney et al., 2022). Also, like LLMs, meta-statistical models could exhibit unexpected failure cases and lack strict guarantees of validity. However, they also offer a promising testbed for mechanistic interpretability (Olah et al., 2020) research: they process structured numerical inputs without tokenization, operate in a single forward pass, and construct mathematical representations rather than linguistically ambiguous ones.